What is the square root of 3x-2=4?
Answer by Fombitz(32388)

sqrt(3x-2)=4
Square both sides.
3x-2=16
3x=18
x=6
Check the answer.
sqrt(3x-2)=4
sqrt(18-2)=4
sqrt(16)=4
4=4
True statement, good answer.